OpenStudy (anonymous):

Rewrite the equation y=x^2-14x+6 in vertex form.

OpenStudy (solomonzelman):

\(\large\color{slate}{y=x^2-14x+6}\) \(\large\color{slate}{y=(x^2-14x)+6}\) \(\large\color{slate}{y=(x^2-14x+49-49)+6}\) \(\large\color{slate}{y=(x^2-14x+49)-49+6}\)

OpenStudy (solomonzelman):

go from there

OpenStudy (anonymous):

Um... what are the next steps?

OpenStudy (solomonzelman):

do you see a perfect square trinomial inside the parenthesis?

OpenStudy (anonymous):

yes?

OpenStudy (solomonzelman):

I didn't change the value of the right side as I wrote the additional `+49-49` in the parenthesis, then I took the `-49` outside the parenthesis. what you need to do, is to factor in the parenthesis

OpenStudy (anonymous):

I am sorry to say I need help with factoring? I missed the class that taught this and not really sure how to complete it.

OpenStudy (solomonzelman):

\(\large\color{slate}{ x^2-14x+49 }\) is same as, \(\large\color{slate}{ x^2-7x-7x+49 }\)

OpenStudy (solomonzelman):

how about now, can you factor it?

OpenStudy (anonymous):

um....

OpenStudy (anonymous):

I work better when someone explains how they got the answer so I can apply it to the other similar problems.

OpenStudy (dtan5457):

Do you still need help on this?

OpenStudy (anonymous):

Yes, thanks!

OpenStudy (dtan5457):

Ok. Just know that vertex form is going to be equal to quadratic formula, and they will equal the same thing, they are just placed in a different way. quadratic: y=ax^2+bx+c vertex: y=a(x-h)^2+k y=x^2-14x+6 complete the square by subtracting the 6 y-6=x^2-14x use (b/2)^2 to get the c value, which would be 49 in this case, add that to both sides again y+43=x^2-14+49 factor that trinomial to (x-7)(x-7) or (x-7)^2 subtract the 43 y=(x-7)^2-43 you can graph both equations and see that they are identical

OpenStudy (dtan5457):

The reason why you must complete the square is because the vertex form of a quadratic requires a perfect square trinomial. That's why you have to subtract and move all those numbers around.

OpenStudy (dtan5457):

Clearer? @djibben615

OpenStudy (anonymous):

gotcha! so the equation y=(x-7)^2-43 is in vertex form and i can graph it now? :) Thanks!

OpenStudy (dtan5457):

Yep! I hope you understand now

OpenStudy (anonymous):

thanks so much!

OpenStudy (dtan5457):

welcome. any more questions?

OpenStudy (anonymous):

Do you know what the value(s) of x is |x| -x = 0? a) x <= 0 b) x = 0 only c) x >= 0 d) no values of x

OpenStudy (anonymous):

@ dtan5457

OpenStudy (dtan5457):

are you missing a > or < by any chance

OpenStudy (anonymous):

nope